OpenStudy (anonymous):

What fraction is the same as 62/8? An answer is required

OpenStudy (anonymous):

62/8 Both the top and bottom are divisible by 2, so we can take a 2 out. 31/4 The numerator and denominator don't share anymore common factors, so the answer is 31/4. That's the fraction in simplest form. Your question only asked for what fraction is the same, and there are an infinite number of answers. Recall that anything multiplied by one is equal to itself. Also recall that anything (except zero) divided by itself is equal to 1. Then, we can say: c/c = 1 c isn't 0 and then x = (c/c) * x If x = 62/8, we can say 62/8 = (c/c) * (62/8) Plug in a c and evaluate. ANY nonzero c will give you an equivalent fraction.

OpenStudy (anonymous):

the answer choices they have are: 7 3/8 7 1/8 Not Here 7 5/8 8 1/4

OpenStudy (anonymous):

62/8 = 7.75 7 3/8 = 7.375 7 1/8 = 7.125 Not Here 7 5/8 = 7.625 8 1/4 = 8.250 It isn't there. And next time include that in your question, lol.

OpenStudy (anonymous):

LOL! I meant to include it but it managed to be left out

OpenStudy (anonymous):

It's 7 6/8, which indeed isn't there (and neither is 7 3/4)